OpenStudy (anonymous):

Solve, cos 2x + 2cosx - 3=0, if 0

OpenStudy (anonymous):

I came up with x=0+2kpi?

OpenStudy (cwrw238):

use the identity cos 2s = 2cos^2 x - 1

OpenStudy (cwrw238):

* cos 2x

OpenStudy (anonymous):

(x+3)(x-1)=0 cos x=-3 cosx=-1 x=0+2kpi

OpenStudy (anonymous):

perhaps I am doing it incorrectly

OpenStudy (cwrw238):

substituting 2cos^2 x - 1 + 2 cos x - 3 = 0 2cos^2 x + 2 cos x - 4 = 0 (2 cos x + 4 )( cos x - 1) = 0 cos x = 1

OpenStudy (anonymous):

so between 0 and 2 pi...

OpenStudy (cwrw238):

well no solution if its if 0<x<2pi but 0 , 2pi if we use if 0=<x<=2pi

OpenStudy (anonymous):

ah, ok, thanks

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

We have cosinuses of 3 values added together ( 2x,x,x) and every single one of them is not bigger then 1 and suming up they are equals to 3( cos2x+2cosx-3=0 <=> cos2x +2cos=3) means they are all equal 1, since if even one of them were lower then one this equasion would make no sense. Thus cos2x=cosx=1, and it is only correct if x=0. Hope i helped. I believe that this reasoning is easier then the one of precursor, as it does not require counting only noticing that cosa is never bigger then 1 no matter the a .